OpenStudy (anonymous):

what is the derivative of 7arcsin(10x^2+15x-9)

OpenStudy (primeralph):

differentiate the stuff inside the arcsin, then multiply that by 7/sqrt(1-(stuff inside arcsin)^2)

OpenStudy (anonymous):

Let \[y=7sin^{-1}(10x^2+15x-9)\] \[\frac{y}{7}=sin^{-1}(10x^2+15x-9)\]\[sin(\frac{y}{7})=(10x^2+15x-9)\]Differentiate both sides with respect to x.

OpenStudy (anonymous):

@_@ i don't see any of those answers as a choice.

zepdrix (zepdrix):

lol he didn't give you the answer, he was giving you a way to solve the problem -_-

zepdrix (zepdrix):

To start, do you know this derivative?\[\large y=\arcsin x \qquad \qquad y'=?\]

OpenStudy (anonymous):

differentiate the stuff inside the arcsin, then multiply that by 7/sqrt(1-(stuff inside arcsin)^2) i did this

OpenStudy (anonymous):

i know it's something like 1/sqrt(1-u)^2

OpenStudy (primeralph):